OpenStudy (sh3lsh):

Find a formula to calculate the sum of the first n odd square terms. How, how do you do this?

OpenStudy (nenadmatematika):

it's quite easy....:D First, I hope you know basic formula for sum of any arithmetic progression Sn=n/2*(a1+an)....where n is the number of elements, a1 is the first element and an is the formula for every element...so 1+3+5+7+....2n-1 is your problem....Sn=n/2*(1+2n-1)=n/2*2n=n^2 :D you can check this formula...for example take sum of the first 4 numbers....1+3+5+7=16 which is 4^2.....

OpenStudy (sh3lsh):

Hahaha. easy?

OpenStudy (nenadmatematika):

ohhh sorry you asked for the sum of square terms 1^2+3^2+...(2n-1)^2...formula is n/3*(2n-1)(2n+1)....sorry if I got you confused :D

OpenStudy (amistre64):

i was wondering what an odd square terms was :)

OpenStudy (nenadmatematika):

:D

OpenStudy (phi):

you could rewrite (2n+1)^2 as 4n^2 + 4n +1 and find the sum of each term.

OpenStudy (phi):

if you follow through on this, you get n(2n-1)(2n+1)/3 where n is the number of terms, starting with 1

OpenStudy (nikvist):